16-Dividends distributed to stockholders should be treated as an expense of the business.

true or false

17-The accrual basis of accounting recognizes revenues when cash is received from customers.

true or false

18-

A company issued 5-year, 10.00% bonds with a par value of $110,000. The market rate when the bonds were issued was 9.50%. The company received $112,315 cash for the bonds. Using the effective interest method, the amount of recorded interest expense for the first semiannual interest period is:

  • $5,500.00.

  • $11,000.00.

  • $5,334.96.

  • $10,607.00.

  • $2,750.00.

  • 21-A company's net sales are $837,340, its costs of goods sold are $456,350, and its net income is $116,370. Its gross margin ratio equals

  • 45.5%.

  • 54.5%.

  • 25.5%.

  • 31.60%.

  • 13.90%.

    24-Required payroll deductions include income taxes, Social Security taxes, pension and health contributions, union dues, and charitable giving.

  • true or false

    27- On December 31 of the current year, the unadjusted trial balance of a company using the percent of receivables method to estimate bad debt included the following: Accounts Receivable, debit balance of $97,800; Allowance for Doubtful Accounts, credit balance of $1,021. What amount should be debited to Bad Debts Expense, assuming 3% of outstanding accounts receivable at the end of the current year are estimated to be uncollectible?

  • $1,021.

  • $1,020.

  • $2,934.

  • $3,955.

  • $1,913.
Answer #1

16. False

Dividend is not a expense for Buisness. It is a part of Appropiration Account to distribute money to investors.

17. False

Accrual Basis of Accounting recognizes the revenue as soon as transaction happen irrespective of whether cash is received or not.

18. $5500

Par Value $110,000

Annual Interest $110,000 * 10% = $11,000

Semiannual interest for first year $11000/2 = $5500

21.  45.5%

Net Sales = $837,340

Cost of Goods Sold = $456,350

Gross Profit = $837340 - $456,350 = $380,990

Gross Margin Ratio = (380,990 / 837,340 ) * 100

= 45.50%

24. False

Required Payroll Deductions are income taxes, Social Security taxes,.

27. $2,934

$97,800 * 3% = $2,934